OpenStudy (anonymous):

please help im not good a math the circumference of a circle is 18n inches. Find its radius and area The radius of the circle is

OpenStudy (anonymous):

the circumference =2*pi*r ,18=2pi*r ,if u devide by 2 pi you get r=8/pi now only you have to subtitute r in the equation of the area which is Area=pi*(r^2)

OpenStudy (anonymous):

what don't you understand on how to solve this problem... do you understand on how to set the problem up? because i think Ahmad 1 is right on his/her explanation

OpenStudy (anonymous):

no i dont understand how to solve it

OpenStudy (anonymous):

ok lets give me one sec

OpenStudy (anonymous):

ok you have 18 inches as the circumference and you are trying to find the radius right? and then by finding the radius you are trying to find the area

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so there are two formulas for circumference: c=2*pi*2 and c=d*pi right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so use the formula c=2*pi*r since you want the radius. substitute c for 18 since you already have that piece of information. you will then set up the formula as 18=2*pi*r ok?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

ok so all that is left to do is to divide 18 by 2 then by pi... because by doing so you will get the radius.... then you can go on to find the area because you have the radius....did i explain it well?

OpenStudy (anonymous):

if you don't get it it is ok it is a lot of information to take in and trying to solve

OpenStudy (anonymous):

im trying to im just not that good at math

OpenStudy (anonymous):

it's ok... i hope i'm helping you not confusing you even more... :)

OpenStudy (anonymous):

no your not confusing me

OpenStudy (anonymous):

ok good cause i was like i hope i'm helping! lol

OpenStudy (anonymous):

so once you divide 18 by 2 then you have the answer

OpenStudy (anonymous):

yes but don't forget to divide it by pi also because you don't want to end up having pi and r on the same side....

OpenStudy (anonymous):

that is what i think if you want to be extra sure than i suggest you to ask your teacher.... but if i'm not mislead then this is the right way..

OpenStudy (anonymous):

I divided it and i got 9 I am not sure if it is right

OpenStudy (anonymous):

did you divide it by pi after dividing it by 2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got 0.5

OpenStudy (anonymous):

ok so did you get the radius to be around 2.86 or 2.87?

OpenStudy (anonymous):

2.86

OpenStudy (anonymous):

ok so you did get 2.86?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh ok that's what i got so now that you have the radius you just plug it into the area formula to get your area and your done! :)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

okay well i hope i helped you! :) good luck!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem! :)